OpenStudy (anonymous):

A motor is pulling a box at a certain rate. You have velocity, a frictional coefficient, and the mass of the box. Find the power needed to keep the box moving at the original velocity and then convert to horsepower..

OpenStudy (anonymous):

This is the same as the previous question you posted. Force of friction is calculated from \[F_f = m \cdot g \cdot \mu\] The motor applies a force equal to the force of friction because the box isn't accelerating. \[P = F_f \cdot v\]